OpenStudy (anonymous):

Express the complex number in trigonometric form. -3i

OpenStudy (freckles):

-3i is the same as 0-3i which is the same as 3(0-1i) so can you find a theta satisfying both of the equations: \[\cos(\theta)=0 \\ \sin(\theta)=-1\]

ganeshie8 (ganeshie8):



ganeshie8 (ganeshie8):

you may use that(unit circle) to find the \(\theta\) corresponding to \((\cos,\sin)= (0,-1)\)